I really need your help in solving this problem. I appreciate your time and efforts. Have a terrific evening :) Here is the question?
A varies directly as z. If A = 30 when z = 5, find A when z = 9

Answer by richard1234(7193)

Just a simple proportion:

30/5 = A/9 --> A = 54

Another way to think of it, A varies directly as z <--> A = kz where k is a constant. Solving 30 = k*5 you get k = 6. When z = 9, A = 6*9 = 54.